OpenStudy (anonymous):

Fan and Metal Please help solve 7(6-2a)=5(-3a+1) and 2r-(5-r)=13+2r

OpenStudy (texaschic101):

7(6 - 2a) = 5(-3a + 1) --- distribute through the parenthesis 42 - 14a = -15a + 5 -- add 15a to both sides 15a - 14a + 42 = 5 -- subtract 42 from both sides 15a - 14a = 5 - 42 -- simplify a = - 37

OpenStudy (texaschic101):

2r - (5 - r) = 13 + 2r -- distribute through the parenthesis 2r - 5 + r = 13 + 2r -- simplify 3r - 5 = 13 + 2r -- subtract 2r from both sides 3r - 2r - 5 = 13 -- add 5 to both sides 3r - 2r = 13 + 5 -- simplify r = 18

OpenStudy (texaschic101):

if you have any questions at all....please ask

OpenStudy (anonymous):

non so far